OVHCloud Web Hosting Status

Current status
Legend
  • Operational
  • Degraded performance
  • Partial Outage
  • Major Outage
  • Under maintenance
FS#11157 — Optical loop Paris
Incident Report for Web Cloud
Resolved
A fibre cable in the customer connection node in Paris has broken. No impact on traffic expected. We will open a ticket with the fibre-optic network operator.

Update(s):

Date: 2014-07-14 11:28:36 UTC
A fiber cable in the in the Robert Keller customer connection node was constrained. The technician intervened and security os once again operational.

Date: 2014-07-14 08:20:52 UTC
The issue has been repaired but there is a loss of 3dB. We will ask the operator to improve the output.

Date: 2014-07-14 05:56:44 UTC
The cut-off was detected 600m from Robert Keller - RKE75.

Because of July 14th,the zone is closed. Reparations can not be performed ​​until Monday evening or Tuesday morning.

Security is operational.

Date: 2014-07-11 09:38:57 UTC
The cut-off is located between the RKE75 and CEV75 sites.
A ticket has been opened with the fibre-optic network operator.
Posted Jul 11, 2014 - 09:37 UTC